OpenStudy (anonymous):

I don't understand how to work this, Please explain and help. What is the axis of symmetry for the function? y = x2 + 3x - 4

OpenStudy (anonymous):

@Mertsj

OpenStudy (compassionate):

The axis of symmetry is given by the equation \[\frac{ -b }{ 2a }\]

OpenStudy (primeralph):

graph it first and look for the lowest point. That vertical line through is the axis

OpenStudy (compassionate):

\[x^2 = a, 3x = b, -4 = c\]

OpenStudy (anonymous):

ok, these are the options I was given... y = -2 y = -1.5 x = -2 x = -1.5

OpenStudy (compassionate):

Okay, so plug it in and do it yourself. \[\frac{ -3 }{ 2(x^2) }\]

OpenStudy (jim766):

it would be \[\frac{ -3 }{2*1 }\]

OpenStudy (compassionate):

\[x^2 = 1\] It is always assumed that x = 1.

OpenStudy (anonymous):

ok...?

OpenStudy (compassionate):

Now do the division.

OpenStudy (anonymous):

not very good at that.... good at polynomials though.

OpenStudy (mertsj):

Do you know that the axis of symmetry is a vertical line that goes through the vertex?

OpenStudy (anonymous):

no...

OpenStudy (mertsj):

Well you do now. Consider this:

OpenStudy (anonymous):

ok

OpenStudy (mertsj):